OpenStudy (anonymous):

How would i do this problem: It says The total number of lights in a triangular rig with n rows is given by the function T(n) as given. Find the number of lights in triangular rigs with 1 to 10 rows. T(n)=1/2n^2+1/2n

OpenStudy (phi):

It sounds like 10 separate problems substitute 1 for n in the formula to get T(1) =0.5 * 1^2 + 0.5*1 = 0.5+ 0.5 = 1 do the same for n=2,3,...10

OpenStudy (anonymous):

oh okay!(:

OpenStudy (phi):

 you could also count them, but that would be tedious